A bus starting from rest moves with a uniform acceleration of 0.1 m s-2 for 2 minutes. Find the distance travelled.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

According to the third equation of motion, v^2- u^2= 2as
s is the distance covered by the bus
(12)^2- (0)^2= 2(0.1) s
s = 720 m
Speed acquired finally by the bus is 12 m/s.
Distance travelled by the bus is 720 m.

Thanks